OpenStudy (anonymous):

MEDAL FOR HELP :) Solve for x: 3(x + 1) = -2(x - 1) + 6.

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

3x + 3 = -2x + 2 + 6 5x = -3 + 2 + 6 5x = 5 x =1

OpenStudy (anonymous):

Can you help me with 1 more? @raimj512

OpenStudy (anonymous):

Yeah, no problem

OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5. @raimj512

OpenStudy (anonymous):

-3x - 9 = -3x - 3 - 5 0x = -17, so there is no x, just something equals 18

OpenStudy (anonymous):

I fanned you

OpenStudy (anonymous):

Why thank you =), by the way, Is that Algebra?